OpenStudy (anonymous):

If Jim Roller can paint a room in 3 hours while John Brush would take 5 hours, how long will they take if they work together? i got 1/8 hour?

OpenStudy (anonymous):

1/3 + 1/5 = 1/x

OpenStudy (anonymous):

x = 15/8 hours

OpenStudy (saifoo.khan):

\[\Huge 1\frac{7}{8} hours.\]